if you know that with 8 units of output, avarage fixed cost is 12.50 dollar and avarage variable cost is 81.25 dollar then what is the total cost at this output level?
1
Expert's answer
2020-11-30T16:42:40-0500

If you know that with 8 units of output, avarage fixed cost is 12.50 dollar and avarage variable cost is 81.25 dollar, then the total cost at this output level is:

"TC = (AFC + AVC)*Q = (12.5 + 81.25)*8 = 750."